OpenStudy (anonymous):

Help plz solve the volume formula for a pyramid, v= 1/3 lwh for h

OpenStudy (anonymous):

@AwesomeSays3!

OpenStudy (anonymous):

whats the problem?

OpenStudy (anonymous):

v= 1/3 *L*W*H use LHS (left hand side)= RHS (right hand side) BY ISOLATING "W" and use opposit of multiply is divided by Hence bring 3 over to LHS which is v = 3*V 3v = LWH W = (3*v )/( L*H) (3 times v then result divided by L and H)

OpenStudy (anonymous):

taa daa

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np